# Mitigating the Impact of Malware Evolution on API Sequence-based Windows Malware Detectors

### v1

@@ Abstract
In dynamic Windows malware detection, deep learning models are extensively deployed to analyze API sequences. Methods based on API sequences play a crucial role in malware prevention. However, due to the continuous updates of APIs and the changes in API sequence calls leading to the constant evolution of malware variants, the detection capability of API sequence-based malware detection models significantly diminishes over time. We observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics. Specifically, compared to the original samples, evolved malware samples often use the API sequences of the pre-evolution samples to achieve similar malicious behaviors. For instance, they access similar sensitive system resources and extend new malicious functions based on the original functionalities. In this paper, we propose a frame(MME), a framework that can enhance existing API sequence-based malware detectors and mitigate the adverse effects of malware evolution. To help detection models capture the similar semantics of these post-evolution API sequences, our framework represents API sequences using API knowledge graphs and system resource encodings and applies contrastive learning to enhance the model’s encoder. Results indicate that, compared to Regular Text-CNN, our framework can significantly reduce the false positive rate by 13.10% and improve the F1-Score by 8.47% on five years of data, achieving the best experimental results. Additionally, evaluations show that our framework can save on the human costs required for model maintenance. We only need 1% of the budget per month to reduce the false positive rate by 11.16% and improve the F1-Score by 6.44%.
Index Terms:
API sequence-based malware detection, Malware evolution, API knowledge graph, Contrastive learning, Deep learning.I Introduction
Malware often executes its malicious activities through a specific sequence of system API calls. Using deep neural networks (DNNs) to analyze and identify these API sequences is proven to be effective in dynamic malware detection [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The research conducted in recent years has focused on achieving high accuracy and minimizing false alarm rates. However, malware detectors are deployed in dynamic environments, where malware variants keep evolving, causing the false negative rate to increase significantly over time [11, 12]. This problem is defined as model aging or concept drift [13]. According to the Kaspersky report in 2019 [14], the false negative rate of a malware detector increased sharply from almost zero to over 20% in just three months. Therefore, mitigating the adverse effects of malware evolution is critical in real malware detection environments.
There appear to be two broad approaches to tackle the malware evolution. The first is to retrain and update detection models with newly labeled samples using online learning [15] or active learning [13], or reject evolved samples until they can be expertly analyzed [16, 17]. However, labeling samples and retraining the model still requires a lot of expert knowledge and computing resources, which incurs a huge cost. The second is to extend the “shelf life” of the model through robust model design and feature space optimization. Malware features are represented to be more robust against temporal bias and reduce the impact of malware evolution [18, 19, 20, 21] However, in existing studies, the raw features extracted mainly focus on the statistical information of static analysis (such as byte histogram, API occurrence, etc), which is not applicable to dynamic detection based on API sequence analysis.
@@
V2 replaces RegOpenKeyEx in V1 with RegOpenKeyTransacted, which means V2 uses transactions to perform the same malicious behavior as V1 for stability. Although the API names of RegOpenKeyEx in V2 and RegOpenKeyTransacted in V1 are different, they represent the same behavior. Intuitively, during evolution, the samples often keep the similar behaviors with different implementations using semantically equivalent APIs.
@@
V2 still uses some APIs that used in V1 (i.e., RegSetValueEx and RegCloseKey). Actually, during evolution, malware samples often involve massive code reuse and generate similar API sequence fragments.
Based on these observations, we proposed to mitigate the impact of malware evolution on API sequence-based windows malware detectors. Our insight is to capture the semantic similarities (including equivalent APIs, similar system resources, and similar API fragments) between the API sequences before and after evolution, and reduce the feature gaps caused by evolution, thus slowing down model aging.
In this paper, we design a framework to Mitigate the impact of Malware Evolution (called MME) to enhance the API sequence-based malware detectors from two perspectives. Specifically, to capture the similarity in API sequences and system resource calls, we have developed a novel API embedding method. The API embedding includes both API name embedding and API parameter embedding. For API name embedding, we analyze the Windows API documents [24] and construct a API knowledge graph, which can capture the similarities between APIs and represent API names as semantic feature vectors. For API argument embedding, the system resources operated by each API are extracted from API arguments and represented as fixed feature vectors. These feature vectors are concatenated and inputed to the detection model.
Second, we enhance the model’s attention to the similar API sequence fragments by designing a contrastive learning strategy. In contrastive learning, the encoder of detection model can measure the similarity of two API sequences by calculating the distance between their two embeddings. Our contrastive learning strategy is to make malware closer to samples of the same family in the feature space and farther away from benign samples. Thus, when a malware sample experiences gradual evolution, it can be expected that new samples will be similar to past samples (as they all have similar API sequence fragments) and hence the contrastive encoder may automatically adapt to evolution.
@@
To evaluate our approach, MME is used to enhance two classic API sequence detection models, namely long short term memory networks (LSTM) and text convolutional neural networks (Text-CNN), as too many work use their variants or combinations as detection models [3, 4, 10, 7, 5, 6]. We collect about 76K Windows PE samples spanning from 2017 to 2021. We train the regular models and enhanced models using data in 2017 and evaluate the performance of them from 2018 to 2021. Our evaluation shows that MME can significantly mitigate the model aging of the malware detectors. It reduces the average false negative rate from 22.4% to 10.1% for LSTM, and from 22.7% to 9.6% for TextCNN. Additionly, MME can significantly reduce the amount of human analyst effort required for model periodical retraining maintenance. The number of samples needed to be labeled can be reduced by 24.19%-94.42%. Finally, model ablation analysis and feature stability analysis explore why MME can help the model mitigate the impact of malware evolution.
@@
We first observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics including equivalent APIs, similar system resources, and similar API fragments. This provides an opportunity to reduce the feature gaps caused by evolution, and slowing down model aging (§I).
@@
We apply MME to two widely used Windows malware detection models. The results show that MME can significantly reduce the high false negtive rete caused by malware evolution, thereby slowing down model aging. MME also can significantly save the human labeling efforts when retraining models (§VII).
@@ II Design Overview
2 Figure 2 shows the overview architecture of MME. Generally, a DNN malware detection model consists of three parts: API sequence embedding, encoder, and classifier. First, API sequence embedding represents each raw API sequence as feature vectors (i.e., embedded API sequence) and input them to the encoder. Then, the encoder learns the features and maps each embedded API sequence to the feature space. Finally, the classifier learns the samples in the feature space and outputs the prediction results (i.e., malware or goodware).
Our framework MME focuses on enhancing the API sequence embedding and encoder modules. For API sequence embedding enhancement, we first construct an API knowledge graph which can find semantically equivalent APIs and using graph embedding to represent API names (§III). Then, to capture the system resources operated by each API, we use feature hash embedding to represent the arguments of each API (§IV). For encoder enhancement, we design a contrastive learning strategy to help model learn the similarities of samples in the same malware family, while learning the dissimilarity between malware and goodware (§V). Finally, the enhancement can be achieved simply by adding MME’s API sequence embedding and contrastive learning strategy to the original model, without altering the original model structure.
@@ III API Name Embedding
We first explain how the konwledge graph can capture the semantic similarity between APIs. The components of API documentation, using RegOpenKeyEx as an example, are shown in Figure 3. Some API-related entities can be extracted from this document, such as action Open (mentioned in the first sentence of the description), prototype RegOpenKey (remove the suffix of the API name), header winreg.h (mentioned in the title), and formal parameters (mentioned in the syntax). Figure 4 shows a small part of the knowledge graph, which captures the relations between the equivalent APIs of RegOpenKeyEx and RegOpenKeyTransacted. Intuitively, these two APIs use the same action, extend from the same prototype, and import from the same header. Besides, they have very similar input/output parameters. That is, these APIs are similar enough in terms of their neighborhoods in the graph. If two APIs are connected to more identical entities, their semantics will become more similar. Therefore, the knowledge graph can capture the similarity between equivalent APIs and then help detectors to detect evolved malware.
@@ III-A1 API Documents Collection
To construct the knowledge graph, the Windows API documents are collected. As shown in Figure 3, each document consists of four parts: title, description, syntax and other information. Among these four parts, title and syntax are structured texts, which contain the basic information of the API (i.e., API name, source header file, class to which it belongs, and function declaration). The description and other information are unstructured texts that contain specific descriptions of API functions and the relationship between the current API and other APIs. We downloads the API documents for Windows 10 from the official website [24] and analyze them to construct an API knowledge graph.
@@ III-A2 Knowledge Graph Construction
For relations, we extracted a total of eight relations (as shown in Table II). Among them, six types of relations can be directly established after entities extraction:
@@
Furthermore, the remaining two types of relations, namely bundled_with and replaced_by, are used to describe the relationships between APIs. The bundled_with refers to the relation that two APIs must be used at the same time, such as a program must call DestroyWindow once for every time it called CreateWindow. The replaced_by means that the two APIs are functionally equivalent and can be used instead, such as RegOpenKeyEx and RegOpenKeyTransacted. These two types of relations can be derived from the unstructured text within API documentation. However, manually extracting these relations one by one from unstructured text is impractical due to the large number of API documents involved. We have observed that there are common patterns when describing the relations between API entities. These patterns can be summarized with templates and utilize them for relation extraction. The template-based relation extraction involves three steps. Firstly, for all API documents, we employ NLP tools to tokenize each unstructured text into sentences and normalize the sentences. Secondly, we select sentences that contain more than one API entity to form a corpus. Thirdly, we employ a semi-automated strategy to analyze the sentences in the corpus and iteratively formulate templates for relation matching. Table III provides several example templates in regular expression format for relations of bundled_with and replaced_by. The detailed process is as follows:
@@
ii) Sentence selection. We employ named entity recognition to extract entities from each sentence, and select sentences that contain more than one API entity to form a corpus. After this step, the scale of the data we need to analyze has been reduced from about 40K API documents to 10K sentences in the cropus.
iii) Template iteratively generation. For each sentence in the corpus, we manually check whether there is bundled_with or replaced_by relation between two API entities. If the answer is no, we remove that sentence from the corpus. Otherwise, we manually formulate a template for the relation and use it for regular expression matching with the all sentences in the corpus. For the sentences that match this template, we extract the corresponding relation from the sentence and remove the sentence from the corpus. Finally, we repeat this process until there is no sentence in the corpus.
@@ IV-A Extract System Resources from Arguments
Figure 5 shows an example hooked API whose name is NtCreateFile. For the first argument, its type is integer, and the value is 2. For the second argument, its type is string and the value is “C:UserAdministratorAppData...” which is a accessed file path.
@@ IV-B Feature Hash Embedding
We use feature hashing [30] to represent each extracted argument as a fixed-length feature vector. Let denotes an substring set of the extracted string argument, and denotes a substring. Let denotes the number of bins. The value of the -th bin is calculated by
@@
The arguments with a large number of shared substrings will have the similar set and will be represented very similar. In this way, if the malware accesses similar system resources before and after evolution, then their feature vectors will be very close.
@@ V Contrastive Encoder
Based on our observations, during evolution, malware samples often involve massive code reuse and generate similar API sequence fragments. In this section, we enhance the encoder’s attention to the similar API sequence fragments by designing a contrastive learning strategy. Through contrastive learning, the encoder can measure the similarity of two embedded API sequences by calculating the distance between them, and make malware closer to samples with similar API fragments and farther away from benign samples in the feature space. Thus, when a malware sample experiences gradual evolution, it can be expected that the representation of new samples will be similar to past samples and the contrastive encoder can automatically adapt to evolution.
As shown in Figure 6, given the input samples with feature vectors, the contrastive learning encoder aims to map them into a latent feature space. Before contrastive learning, the evolved malware produce many differences in the feature space, leading the detection model to misclassify it as benign. Then, the contrastive learning optimizes the encoder and generates a latent space. In the latent space, pairs of samples in the same class have a smaller distance, and pairs of samples from different classes have a larger distance. As such, the encoder will pay more attention to the similarities among samples from the same malware family. Any evolved sample that retains similar API fragments to the past samples will be represented as closer, thereby reducing misclassifications of the evolved samples.

### v2

@@ Abstract
In dynamic Windows malware detection, deep learning models are extensively deployed to analyze API sequences. Methods based on API sequences play a crucial role in malware prevention. However, due to the continuous updates of APIs and the changes in API sequence calls leading to the constant evolution of malware variants, the detection capability of API sequence-based malware detection models significantly diminishes over time. We observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics. Specifically, compared to the original samples, evolved malware samples often use the API sequences of the pre-evolution samples to achieve similar malicious behaviors. For instance, they access similar sensitive system resources and extend new malicious functions based on the original functionalities. In this paper, we propose a framework MME(Mitigating the impact of Malware Evolution), a framework that can enhance existing API sequence-based malware detectors and mitigate the adverse effects of malware evolution. To help detection models capture the similar semantics of these post-evolution API sequences, our framework represents API sequences using API knowledge graphs and system resource encodings and applies contrastive learning to enhance the model’s encoder. Results indicate that, compared to regular Text-CNN, our framework can significantly reduce the false positive rate by 13.10% and improve the F1-Score by 8.47% on five years of data, achieving the best experimental results. Additionally, evaluations show that our framework can save on the human costs required for model maintenance. We only need 1% of the budget per month to reduce the false positive rate by 11.16% and improve the F1-Score by 6.44%.
I Introduction
Malware typically executes its malicious activities through a series of specific system API calls. These API call sequences not only reflect the behavioral patterns of malware but also provide critical dynamic features for detection systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Using deep neural networks (DNNs) to analyze and identify these API sequences has proven highly effective in dynamic malware detection, as they can capture temporal dependencies and complex patterns within the sequences. However, with the continuous evolution of malware, detection models face challenges such as model aging or concept drift [11, 12], which lead to a significant increase in false positive rates over time [13]. Therefore, in-depth research on API sequence analysis to mitigate the impact of malware evolution is crucial for building robust detection systems.
Existing research has proposed various methods to tackle the challenges posed by malware evolution. One common strategy involves retraining models with newly labeled samples through online learning [14] or active learning [13][15], or rejecting evolved samples for expert analysis [16, 17, 18]. However, these approaches demand substantial expert knowledge and computational resources, rendering them costly. An alternative strategy focuses on extending model lifespan by optimizing feature spaces and designing robust models [19, 20, 21, 22]. For instance, APIGraph [21] constructs an API knowledge graph to represent API feature vectors and replaces original inputs with clustered API identifiers, ensuring that APIs with similar semantics share identical cluster identifiers. This method emphasizes semantic similarity between APIs and reduces feature space dimensionality through clustering. Similarly, SDAC [22] calculates semantic distances between APIs based on differences in API vectors within sequences, clustering APIs according to these distances and focusing on their semantic relationships within sequences. While these methods have achieved some success in capturing API-level similarities, they exhibit limitations: they fail to fully exploit system resource access patterns within API parameters and offer limited capability to intervene in existing models.
@@ I Introduction
V2 replaces RegOpenKeyEx in V1 with RegOpenKeyTransacted, which means V2 uses transactions to perform the same malicious behavior as V1 for stability. Although the API names of RegOpenKeyEx in V2 and RegOpenKeyTransacted in V1 are different, they represent the same behavior. Intuitively, during evolution, the samples often keep similar behaviors with different implementations using semantically equivalent APIs.
@@
V2 still uses some APIs that used in V1 (i.e., RegSetValueEx and RegCloseKey). Actually, during evolution, malware samples often involve massive code reuse and generate similar fragments of the API sequence.
Based on these observations, in this paper, we propose a framework named MME(Mitigating the impact of Malware Evolution) designed to mitigate model aging and enhance API sequence-based malware detectors from four key perspectives.
(1) Mining API Parameter Features. MME leverages dynamic analysis to acquire API calls along with their parameters. It employs hierarchical hash encoding on these parameters to extract system resource access patterns (e.g., file paths and registry keys), thereby providing a more comprehensive representation of malware behavior. By fully exploiting the information embedded within API parameters, MME improves the capture of malicious activities. For parameter embedding, system resources accessed by each API operation are extracted from the parameters and represented as fixed feature vectors. These vectors are subsequently concatenated and fed into the detection model.
(2) Enhancing the Model Encoder. MME incorporates contrastive learning to strengthen the model’s encoder, boosting its generalization capability against evolving malware. By integrating knowledge graph embeddings with contrastive learning objectives, the encoder measures the similarity between two API sequences by computing the distance between their embedding vectors. Our contrastive learning strategy ensures that malware samples in the feature space are closer to other samples within the same family while maintaining a distinct separation from benign samples. Consequently, as malware undergoes gradual evolution, new samples are expected to resemble past samples (due to shared API sequence fragments), enabling the contrastive encoder to adapt automatically to this evolution. This approach enhances the encoder’s ability to capture both semantic and temporal robustness.
(3) Constructing a Knowledge Graph for Richer Semantic Representation. MME constructs a knowledge graph using API documentation [25] to capture semantic relationships between APIs. This construction process is more thorough, integrating API function descriptions and parameter information to deliver a richer semantic representation. As a result, MME effectively identifies similarities between APIs, enhancing the overall understanding of their behavioral context.
(4) Augmenting Existing Models. MME is designed to enhance existing API sequence-based models, improving their robustness in dynamic environments. For instance, it can be seamlessly integrated with models such as LSTM [26], Text-CNN [27], Transformer [28], and BERT [29], further elevating their performance. This adaptability makes MME a versatile solution for advancing malware detection capabilities.
@@
To evaluate our approach, MME is used to enhance two classic API sequence detection models, namely long short-term memory networks (LSTM) and text convolutional neural networks (Text-CNN), as many work use their variants or combinations as detection models [3, 4, 10, 7, 5, 6]. We collect about 76K Windows PE samples spanning from 2017 to 2021. We train the regular models and enhanced models using data in 2017 and evaluate their performance of them from 2018 to 2021. Our evaluation shows that MME can significantly mitigate the model aging of the malware detectors. It reduces the average false negative rate from 22.4% to 10.1% for LSTM, and from 22.7% to 9.6% for TextCNN. Additionly, MME can significantly reduce the amount of human analyst effort required for model periodical retraining and maintenance. The number of samples needed to be labeled can be reduced by 24.19%-94.42%. Finally, model ablation analysis and feature stability analysis explore why MME can help the model mitigate the impact of malware evolution.
@@
We first observe that the API sequences of malware samples before and after evolution usually have similar malicious semantics, including equivalent APIs, similar system resources, and similar API fragments. This provides an opportunity to reduce the feature gaps caused by evolution, and to slow down model aging (§I).
@@
We apply MME to two widely used Windows malware detection models. The results show that MME can significantly reduce the high false negative rate caused by malware evolution, thereby slowing down model aging. Thus, the aging speed of the model is slowed down. At the same time, MME can also significantly save the workload of manual labeling when retraining the model, showing its universality and the double optimization of reducing labor costs. (§VII).
@@ II Design Overview
Figure 2 shows the overview architecture of MME. Generally, a DNN malware detection model consists of three parts: API sequence embedding, encoder, and classifier. First, API sequence embedding represents each raw API sequence as feature vectors (i.e., embedded API sequence) and inputs them to the encoder. Then, the encoder learns the features and maps each embedded API sequence to the feature space. Finally, the classifier learns the samples in the feature space and outputs the prediction results (i.e., malware or goodware).
Our framework MME focuses on enhancing the API sequence embedding and encoder modules. For API sequence embedding enhancement, we first construct an API knowledge graph that can find semantically equivalent APIs and use graph embedding to represent API names (§III). Then, to capture the system resources operated by each API, we use feature hash embedding to represent the arguments of each API (§IV). For encoder enhancement, we design a contrastive learning strategy to help the model learn the similarities of samples in the same malware family, while learning the dissimilarity between malware and goodware (§V). Finally, the enhancement can be achieved simply by adding MME’s API sequence embedding and contrastive learning strategy to the original model, without altering the original model structure.
@@ III API Name Embedding
We first explain how the knowledge graph can capture the semantic similarity between APIs. The components of API documentation, using RegOpenKeyEx as an example, are shown in Figure 3. Some API-related entities can be extracted from this document, such as action Open (mentioned in the first sentence of the description), prototype RegOpenKey (remove the suffix of the API name), header winreg.h (mentioned in the title), and formal parameters (mentioned in the syntax). Figure 4 shows a small part of the knowledge graph, which captures the relations between the equivalent APIs of RegOpenKeyEx and RegOpenKeyTransacted. Intuitively, these two APIs use the same action, extend from the same prototype, and import from the same header. Besides, they have very similar input/output parameters. That is, these APIs are similar enough in terms of their neighborhoods in the graph. If two APIs are connected to more identical entities, their semantics will become more similar. Therefore, the knowledge graph can capture the similarity between equivalent APIs and then help detectors to detect evolved malware.
@@ III-A1 API Documents Collection
To construct the knowledge graph, the Windows API documents are collected. As shown in Figure 3, each document consists of four parts: title, description, syntax and other information. Among these four parts, title and syntax are structured texts, which contain the basic information of the API (i.e., API name, source header file, class to which it belongs, and function declaration). The description and other information are unstructured texts that contain specific descriptions of API functions and the relationship between the current API and other APIs. We downloaded the API documents for Windows 10 from the official website [25] and analyzed them to construct an API knowledge graph.
@@ III-A2 Knowledge Graph Construction
For relations, we extracted a total of eight relations (as shown in Table II). Among them, six types of relations can be directly established after entity extraction:
@@
Furthermore, the remaining two types of relations, namely bundled_with and replaced_by, are used to describe the relationships between APIs. The bundled_with refers to the relation that two APIs must be used at the same time, such as a program must call DestroyWindow once for every time it called CreateWindow. The replaced_by means that the two APIs are functionally equivalent and can be used instead, such as RegOpenKeyEx and RegOpenKeyTransacted. These two types of relations can be derived from the unstructured text within the API documentation. However, manually extracting these relations one by one from unstructured text is impractical due to the large number of API documents involved. We have observed that there are common patterns when describing the relations between API entities. These patterns can be summarized with templates and used them for relation extraction. The template-based relation extraction involves three steps. Firstly, for all API documents, we employ NLP tools to tokenize each unstructured text into sentences and normalize the sentences. Secondly, we select sentences that contain more than one API entity to form a corpus. Thirdly, we employ a semi-automated strategy to analyze the sentences in the corpus and iteratively formulate templates for relation matching. Table III provides several example templates in regular expression format for relations of bundled_with and replaced_by. The detailed process is as follows:
@@
ii) Sentence selection. We employ named entity recognition to extract entities from each sentence, and select sentences that contain more than one API entity to form a corpus. After this step, the scale of the data we need to analyze has been reduced from about 40K API documents to 10K sentences in the corpus.
iii) Template iteratively generation. For each sentence in the corpus, we manually check whether there is bundled_with or replaced_by relation between two API entities. If the answer is no, we remove that sentence from the corpus. Otherwise, we manually formulate a template for the relation and use it for regular expression matching with all the sentences in the corpus. For the sentences that match this template, we extract the corresponding relation from the sentence and remove the sentence from the corpus. Finally, we repeat this process until there are no sentences in the corpus.
@@ IV-A Extract System Resources from Arguments
Figure 5 shows an example hooked API whose name is NtCreateFile. For the first argument, its type is integer, and the value is 2. For the second argument, its type is string and the value is “C:UserAdministratorAppData...” which is an accessed file path.
@@ IV-B Feature Hash Embedding
We use feature hashing [35] to represent each extracted argument as a fixed-length feature vector. Let denotes a substring set of the extracted string argument, and denotes a substring. Let denote the number of bins. The value of the -th bin is calculated by
@@
The arguments with a large number of shared substrings will have the similar set and will be represented very similarly. In this way, if the malware accesses similar system resources before and after evolution, then their feature vectors will be very close.
@@ V Contrastive Encoder
Based on our observations, during evolution, malware samples often involve massive code reuse and generate similar API sequence fragments. In this section, we enhance the encoder’s attention to similar API sequence fragments by designing a contrastive learning strategy. Through contrastive learning, the encoder can measure the similarity of two embedded API sequences by calculating the distance between them, and make malware closer to samples with similar API fragments and farther away from benign samples in the feature space. Thus, when a malware sample experiences gradual evolution, it can be expected that the representation of new samples will be similar to past samples and the contrastive encoder can automatically adapt to evolution. Our approach focuses on family-level similarity and seeks to ensure that malware variants evolving within the same family are closely embedded in the feature space, thus mitigating model aging caused by API substitutions and sequence variations.
As shown in Figure 6, given the input samples with feature vectors, the contrastive learning encoder aims to map them into a latent feature space. Before contrastive learning, the evolved malware produces many differences in the feature space, leading the detection model to misclassify it as benign. Then, the contrastive learning optimizes the encoder and generates a latent space. In the latent space, pairs of samples in the same class have a smaller distance, and pairs of samples from different classes have a larger distance. As such, the encoder will pay more attention to the similarities among samples from the same malware family. Any evolved sample that retains similar API fragments to the past samples will be represented as closer, thereby reducing misclassifications of the evolved samples.